OpenStudy (anonymous):

Hunter placed 2 red, 3 blue, 1 yellow, 8 black, and 7 white straws into a bag what is the probability that he will pull out a white straw whwn he reaches into the bag

OpenStudy (anonymous):

there are a total of 2+3+1+8+7=21 straws. There is only 7 white straws. Thus P(white)=7/21 = 1/3